# Fast and Private Inference of Deep Neural Networks by Co-designing Activation Functions

### v1

@@ 5.1 The Problem with Polynomial Activation Functions
The first step in replacing an activation function with a polynomial is to design a polynomial that approximates the original function as closely as possible. A common approach for this is the least-squares polynomial fitting. In this approach, a small discretized range is chosen to fit the polynomial on. A table of values is created for the function over all values in the range. This creates a system of equations for the polynomial coefficients that can be solved with least squares. The challenge with this approach is that, outside of this range, the polynomial no longer resembles the original activation function and often diverges rapidly. This leads to a problem called escaping activations, first identified by Garimella et al. [12]. If one naively swaps a ReLU for its polynomial approximation, with no additional modifications to the training, all weights will become infinite within a few training epochs. We give an example of this degradation in Figure 5. We can see that, without modifying the training procedure, a polynomial can completely destroy the accuracy.
To illustrate the problem more clearly, we conduct an experiment using a polynomial of degree four fitted on the range $[-5,5]$ ($\lambda$=5) as the activation function for a three-layer model on CIFAR-10 [21]. In Figure 5, we plot this polynomial activation function and the $\ell_{\infty}$-norm of the input and output to each activation function. We note that, with no modification (except replacing ReLUs with polynomials), the weights of this model become undefined within approximately three epochs of training. First, we note the divergent behaviour of the polynomial outside the fitted range. Second, we observe the effect of the divergence on the outputs of the activation function. Specifically, we wait until the model weights become undefined (NaN in Python) and then observe the behaviour leading up to the explosion. We can see that three steps before the model weights become undefined (NaN), the input values of each activation are out-of-range, but the outputs still behave similarly to a ReLU. However, in the next iteration (two steps before NaN), a single value in the first layer goes too far out of range. This causes a ripple effect for the other two layers, creating an extremely large output (approx $-2000$) in the final activation function. This large value creates a large gradient, and after another iteration of training, the values become so large that the gradients (and weights) become undefined (NaN). We find that training the model by minimizing the classification loss alone is not enough to keep the model in range as the gradients explode before decreasing the loss.
@@
An additional challenge is that we will evaluate the fitted polynomial in a finite ring with limited precision. This significantly impacts the polynomial coefficients, which tend to be relatively small, especially for the higher-order terms. Specifically, these small coefficients can get truncated to zero in limited precision, which causes the polynomial to diverge even inside the fitted range. We give an example of this in Appendix B.
@@ 5.2 Defining PILLAR
Next, we solve the system $AB=Y$ for $A$ using mixed integer linear programming with $A\in[-2^{p}-1,2^{p}-1]$ to get the coefficients $A$ that minimize the error between the polynomial $AB$ and the ReLU values $Y$. Finally, we scale the resulting coefficients down by $2^{p}$. We note that $A\in[-2^{p}-1,2^{p}-1]$ corresponds to coefficients being bounded by $[-1,1]$ after we scale down. We use 10 bits of precision for all polynomials motivated by our derivations in Section 4. As we observe in Appendix B, our quantized polynomial fitting addresses the problems of exploding activations within the range. However, the issue of going out-of-range requires additional treatment.
@@
Although activation regularization teaches the model not to go out-of-range over time, the model still needs to avoid going to infinity during the early stages of training. Thus, during training, we apply a clipping function to the input of the activation function such that if any input goes out of range, it is truncated to the range’s maximum (or minimum) value. This clipping function does not affect the penalty as it is applied after the penalty function has been computed. We emphasize that this clipping function is only used during training and is removed during inference. The intuition is that the model should learn not to go out-of-range during training and thus no longer requires this clipping function during inference. Additionally, we often find that setting the $\lambda_{reg}$ of the penalty to be smaller than the range used for clipping and polynomial fitting can yield even better results as it gives room for going slightly out-of-range in the inference.
Dataset Model Accuracy $\pm$ CI
Cifar10 MiniONN 88.1 $\pm$ 0.26
VGG 16 90.8 $\pm$ 0.11
ResNet18 93.4 $\pm$ 0.14
ResNet110 91.4 $\pm$ 0.18
CIFAR-100 VGG 16 66.3 $\pm$ 0.22
ResNet32 67.8 $\pm$ 0.32
ResNet18 74.9 $\pm$ 0.14
ImageNet ResNet50 77.7
@@ 5.3 Measuring PILLAR’s Effectiveness
We evaluate PILLAR across a range of different models and architectures considered in related work [17, 29, 32]. We summarize the results in Table 1. All results are averaged over five random seeds, and we show the $95\%$ confidence interval. The only exception is ResNet50 on ImageNet, where we only train a single model due to the size of the dataset. We defer to Section 6 for the details of the experimental setup. We note that these results are using PyTorch with no cryptography or quantization. We give a complete evaluation using MPC in Section 6 where quantization has an effect. Table 1 provides preliminary evidence that our polynomial training approach yields high accuracies competitive with state-of-the-art ReLU models across a range of models and datasets.
@@ 6 Evaluation of Co-Design
[scale=0.9] {axis}[ title=, xlabel=Inference Time (s), ylabel=Test Accuracy (%), xmode=log, grid=major, ymin=60, ymax=80, legend pos=south west, ]
[…]
@@
In this section, we provide an end-to-end comparison of our co-design against state-of-the-art solutions in secure inference. We determine the state-of-the-art works following a recent SoK by Ng and Chow [30]. Specifically, we consider three solutions on the Pareto front of latency and accuracy as determined by Ng and Chow. These works are COINN [17], GForce [29] and CrypTen (CryptGPU) [20, 32]. We will evaluate the metrics of latency (or runtime of a single sample) and encrypted accuracy. We begin with the experimental setup, then evaluate both metrics against each related work. Section 6.2 evaluates the ResNet-18 architecture, which gives our state-of-the-art performance. Section 6.3 evaluates the VGG-16 architecture, the only architecture GForce evaluates. Section 6.4 considers other ResNets and the MiniONN architecture following COINN. Finally, in Section 6.5, we evaluate ImageNet.
@@
We plot a summary of the accuracy and inference time for CIFAR-100 in Figure 6. For both datasets (recall Figure 1 for CIFAR-10), we observe that our work always gives the solution with the fastest inference time by a statistically significant amount. In terms of accuracy our work is competitive with the state-of-the-art but CryptGPU is always the most accurate as it can infer unmodified plaintext models. We find that our solution is faster than CryptGPU by 4×4$\times$4 ×, GForce by 5×5$\times$5 ×, and COINN by 18×18$\times$18 × on average in wide area networks. Our accuracies are competitive with state-of-the-art and plaintext solutions and stay stable (no escaping activations) with models containing up to $110$ layers and $23$ million parameters.
@@ 6.1 Experimental Setup
We develop an experimental setup that follows as closely as possible to the works we compare to [17, 29, 32]. We use CIFAR-10/100 [21] and ImageNet [10], the same common benchmark datasets as related work. Our model architectures include: MiniONN [24], VGG [31], and ResNets [16]. This covers models of depth $7$ to $110$ layers with the number of trainable parameters ranging from $0.2$ to $23$ million. See Appendix C for further details on architectures and datasets.
@@
We implement PILLAR using GEKKO’s [3] mixed integer linear programming solver for the quantization-aware polynomial fitting. We then cache these polynomial coefficients and use them for all models and datasets. We train all our models using PyTorch and implement a custom activation function module for the polynomial approximations. We call this the PolyReLU layer and use it to replace all ReLU layers. This module takes the cached coefficients and computes the output of the polynomial approximation for forward passes. When training, we compute the regularization penalty by taking a snapshot of the inputs passed to the PolyReLU layers and computing our modified loss function from equation 7. After training, we export the model as an ONNX model which can be inputted to CrypTen.
We implement ESPN and HoneyBadger in the CrypTen interface. We add configuration parameters that allow the user to specify which type of activation function they would like from ReLU or PolyReLU. Similarly, the polynomial evaluation method is parametrized in the config file between ESPN, HoneyBadger, and the default CrypTen Polynomial Evaluation. We follow CrypTen’s default trusted first party provider, which assumes a first party will generate and distribute all needed pre-computed values from Beaver triplets to binary shares, etc. To ensure we accurately measure the online phase of the inference, we move some additional operations to the pre-computation phase. Specifically, random-number generation in the Pseudo-Random-Zero-Sharing (PRZs) are fixed to zero to simulate being computed in the offline phase.
For ESPN, we implement the idea directly into CrypTen’s polynomial evaluation module. As disccused in Section 4, we compute the powers for the last term in the polynomial first, and then reuse these powers for all lower-degree terms. For HoneyBadger, we modify the Trusted First-Party Provider to provide the additional sets of random numbers and exponents needed for their protocol. We then implement the idea directly into the polynomial module of ArithmeticSharedTensors. We use HoneyBadger’s proposed dynamic programming method with the memory optimization technique of keeping only the current and previous iterations in memory.
When running our experiments on the GPU, we observed overflows not present in the CPU version of CrypTen. Upon inspection, we found that the default number of blocks ($4$) set in CryptGPU is not adequate for our use-case. Increasing this parameter to $5$ fixed the overflow issue completely. All experiments are run on a machine with 32 CPU cores @ 3.7 GHz and 1 TB of RAM with two NVIDIA A100 with 80 GB of memory. We simulate network delay by calling the sleep function for the appropriate time whenever the client and server communicate. We simulate the LAN with 0.25 ms roundtrip delay and the WAN with 100 ms, following COINN [17]. All experiments (except ImageNet) are repeated over multiple random seeds, and we report the mean and 95% confidence interval as shaded areas.
Since COINN does not include source code, only executables, we use the numbers reported in the paper for their work. For all other works, we run our own benchmarks. To evaluate GForce we use their code unmodified. In order to use CryptGPU in practice, one must first train a model in PyTorch. For ImageNet, PyTorch provides pre-trained models that we can use. However, we will need to train a model for all other architectures and datasets. We simply use the same configurations as our PolyRelu models but with standard ReLUs. We include all source code to reproduce our results [^1] https://github.com/LucasFenaux/PILLAR-ESPN.
@@
We introduce five new hyperparameters associated with our techniques: polynomial degree ($n$), polynomial approximation range ($\lambda$), polynomial regularization range ($\lambda_{reg}$), polynomial regularization coefficient ($\beta$), and polynomial regularization exponent ($\gamma$). We found that, for all models and datasets, a value of $\gamma=10$ performs well as it introduces a strong enough incentive for PolyReLU inputs to stay in range while keeping penalization for values in range practically 00 (if an input $x$ is within range, then $\frac{x}{\lambda_{reg}}<1\Rightarrow(\frac{x}{(\lambda_{reg})^{10}}\sim 0$). We also found that a polynomial approximation range $\lambda=5$ provides a good compromise between regularization (larger ranges require less regularization) and quantization (lower range values use less precision). Similarly, we found the optimal quantization-aware polynomial degree ($n$) to be 4. This value provides a good trade-off between approximation quality while avoiding overflow by keeping the total precision less than $64$-bits. We vary the polynomial regularization range ($\lambda_{reg}$) and polynomial regularization coefficient ($\beta$) per model and dataset, although we found $\lambda_{reg}=4.8$ (slightly tighter than $\lambda=5$) and $\beta=5\times 10^{-5}$ to be good default values. We use a precision of $p=12$ as this allows us to work in CrypTen’s $64$-bit ring.
We used Stochastic Gradient Decent as the optimizer with a learning rate of $0.013$ as the default. This includes a Cosine Annealing Learning Rate Scheduler with Linear Learning Rate Warmup of 5 epochs and decay $0.01$. We use a weight decay of either $10^{-4}$ or $5\times 10^{-4}$ and a momentum of $0.9$. We used a default batch size of $128$ and set the default number of Epochs to $185$. For some models we tuned the learning rate, number of epochs, and regularization coefficient to achieve a slightly higher accuracy. We detail hyperparameters in our source code repository.
@@ 6.3 VGG-16 Architecture
We recall that we swap the MaxPool layers for AvgPool layers in the inference time evaluation. We note that this comes at a cost to accuracy for the VGG architecture. Thus, to give the best scenario possible for GForce, we consider the accuracy of GForce with MaxPools and our work with AvgPool. We give the results in Table 3. As expected, GForce outperforms our work in accuracy (due to the MaxPools); however, only by a few percentage points. We train CryptGPU to also use AvgPool and find it also loses a few percentage points, confirming our hypothesis that a MaxPool is necessary for high accuracy in the VGG-16 architecture. We emphasize that our ResNet-18 result outperforms GForce’s VGG result in both inference time and accuracy. Furthermore, ResNets are a more popular and compact architecture due to skip-connections.
@@ 6.4 Other Architectures
While GForce is the current state-of-the-art, COINN is a competitive solution that evaluates ResNet architectures. Thus, we also evaluate the same configurations as COINN. This includes the smaller MiniONN architecture, a ResNet-32, and a ResNet-110. We recall that we use the results from the COINN paper directly for this evaluation. We exclude GForce from this evaluation as they only evaluate VGG-16 models.
@@
We once again swap all MaxPool layers for AvgPool in our work and CryptGPU but leave COINN unmodified. We give the results in Figure 9. We observe that, over each of the three increasingly large architectures, the trends are similar and proportional to the number of parameters (0.2, 0.5, and 1.7 million parameters for MiniONN, ResNet-32, and ResNet-110, respectively). Across all architectures and network delays, our work outperforms COINN by a statistically significant amount (18×18$\times$18 × on average in WAN). We once again outperform CryptGPU in all evaluations [^2] All of which are statistically significant except MiniONN in LAN where the confidence intervals overlap slightly. with a 4×4$\times$4 × speed up on average in the WAN.
@@
CIFAR-10 / MiniONN PILLAR 88.1 $\pm$ 0.26 87.85 $\pm$ 0.36
@@
CIFAR-10 / ResNet-110 PILLAR 91.4 $\pm$ 0.18 91.4 $\pm$ 0.18
@@
CIFAR-100 / ResNet-32 PILLAR 67.8 $\pm$ 0.32 67.84 $\pm$ 0.35

### v2

@@ 5.1 The Problem with Polynomial Activation Functions
The first step in replacing an activation function with a polynomial is to design a polynomial that closely approximates the original function. A common approach for this is the least-squares polynomial fitting. In this approach, a table of values is created for the polynomial over a small discretized range of values. This creates a system of equations for the polynomial coefficients that can be solved with least squares. The challenge with this approach is that, outside of this range, the polynomial no longer resembles the original activation function and often diverges rapidly. This leads to a problem called escaping activations, first identified by Garimella et al. [12]. If one naively swaps a ReLU for its polynomial approximation, all weights will become infinite within a few training epochs. We give an example of this degradation in Figure 5. We can see that, without modifying the training procedure, a polynomial can completely destroy the accuracy.
To illustrate the problem more clearly, we conduct an experiment using a polynomial of degree four fitted on the range $[-5,5]$ ($\lambda$=5) as the activation function for a three-layer model on CIFAR-10 [23]. In Figure 5, we plot this polynomial activation function and the $\ell_{\infty}$-norm of the input and output to each activation function. We note that, with no modification (except replacing ReLUs with polynomials), the weights of this model become undefined within approximately three epochs of training. First, we note the divergent behaviour of the polynomial outside the fitted range. Second, we observe the effect of the divergence on the outputs of the activation function. Specifically, we wait until the model weights become undefined (NaN in Python) and then observe the behaviour leading up to the explosion. We can see that three steps before the model weights become undefined (NaN), the input values of each activation are out-of-range, but the outputs still behave similarly to a ReLU. However, in the next iteration (two steps before NaN), a single value in the first layer goes too far out of range. This causes a ripple effect for the other two layers, creating an extremely large output (approx $-2000$) in the final activation function. This large value creates a large gradient, and after another iteration of training, the values become so large that the gradients (and weights) become undefined (NaN). We find that minimizing the classification loss alone is not enough to keep the model in range as the gradients explode before decreasing the loss.
@@
An additional challenge is that we will evaluate the fitted polynomial in a finite ring with limited precision. This significantly impacts the polynomial coefficients, which tend to be relatively small, especially for the higher-order terms. Specifically, these small coefficients can get truncated to zero in limited precision, which causes the polynomial to diverge even inside the fitted range. We give an example of this in Appendix C.
@@ 5.2 Defining PILLAR
Next, we solve the system $AB=Y$ for $A$ using mixed integer linear programming with $A\in[-2^{p}-1,2^{p}-1]$ to get the coefficients $A$ that minimize the error between the polynomial $AB$ and the ReLU values $Y$. Finally, we scale the resulting coefficients down by $2^{p}$. We note that $A\in[-2^{p}-1,2^{p}-1]$ corresponds to coefficients being bounded by $[-1,1]$ after we scale down. We empirically choose $p=10$ for all polynomials during the ML training. As we observe in Appendix C, our quantized polynomial fitting addresses the problems of exploding activations within the range. However, the issue of going out-of-range requires additional treatment.
@@
Although activation regularization teaches the model not to go out-of-range over time, the model still needs to avoid going to infinity during the early stages of training. Thus, during training, we apply a clipping function to the input of the activation function such that if any input goes out of range, it is truncated to the range’s maximum (or minimum) value. This clipping function does not affect the penalty as it is applied after the penalty function has been computed. We emphasize that this clipping function is only used during training and is removed during inference. The intuition is that the model should learn not to go out-of-range during training and thus no longer requires this clipping function during inference. Additionally, we find that setting the $\lambda_{reg}$ of the penalty to be smaller than the range used for polynomial fitting (and clipping) can yield even better results. This is because the polynomial will be accurate for a larger range outside of the range the model was regularized to stay inside, allowing an extra buffer in case of failure during inference.
Dataset Model Plain Accuracy $\pm$ CI
ReLU PILLAR
Cifar10 MiniONN 91.2 $\pm$ 0.17 88.1 $\pm$ 0.26
VGG 16 92.6 $\pm$ 0.16 90.8 $\pm$ 0.11
ResNet18 94.7 $\pm$ 0.09 93.4 $\pm$ 0.14
ResNet110 92.8 $\pm$ 0.27 91.4 $\pm$ 0.18
CIFAR-100 VGG 16 70.9 $\pm$ 0.17 66.3 $\pm$ 0.22
ResNet32 68.4 $\pm$ 0.46 67.8 $\pm$ 0.32
ResNet18 76.6 $\pm$ 0.07 74.9 $\pm$ 0.14
ImageNet ResNet50 80.8 77.7
@@
Summary of PILLAR
We refer to PILLARas the combination of all components described in this section. We note that the clipping and regularization warm-up components are used (only during training) to enable regularization by preventing the model from going to infinity. The regularization component ensures that the model will stay in range during inference. All three of these components play a crucial role in the success of PILLAR, as removing any of them will result in a model with unacceptable accuracy (due to infinite weights or escaping activations). The batch norm is the only optional component, which we give a small ablation study over in Appendix A.
@@ 5.3 Measuring PILLAR’s Effectiveness
We evaluate PILLAR across a range of different models and architectures considered in related work [18, 31, 35]. We summarize the results in Table 1. We include the accuracy of a model trained with standard ReLUs as a baseline. All results are averaged over five random seeds, and we show the $95\%$ confidence interval. The only exception is ResNet50 on ImageNet, where we only train a single model due to the size of the dataset. We defer to Section 6 for the details of the experimental setup. We note that these results are using PyTorch with no cryptography or quantization. We give a complete evaluation using MPC in Section 6 where quantization has an effect. Table 1 provides preliminary evidence that our polynomial training approach yields high accuracies competitive with state-of-the-art ReLU models across a range of models and datasets.
@@ 6 Evaluation of Co-Design
$10^{0.5}$$10^{1}$$10^{1.5}$6060606065656565707070707575757580808080Inference Time (s)Test Accuracy (%)This WorkCOINN [18]GForce [31]CryptGPU [35]
@@
In this section, we provide an end-to-end comparison of our co-design against state-of-the-art solutions in secure inference. We evaluate the performance of PILLAR and Algorithm 2 using both ESPN and HoneyBadger  as they offer different trade-offs in the type of pre-computation needed and the size of the communication. We determine the state-of-the-art works following a recent SoK by Ng and Chow [32]. Specifically, we consider three solutions on the Pareto front of latency and accuracy as determined by Ng and Chow. These works are COINN [18], GForce [31] and CrypTen (CryptGPU) [22, 35]. We also evaluate Cheetah [17], a recent work not included in the SoK. We will evaluate the metrics of latency (or runtime of a single sample) and encrypted accuracy. We additionally evaluate the communication and number of rounds in Appendix D. We begin with the experimental setup, then evaluate both metrics (runtime and accuracy) against each related work. Section 6.2 evaluates the ResNet-18 architecture, which gives our state-of-the-art performance. Section 6.3 evaluates the VGG-16 architecture, the only architecture GForce evaluates. Section 6.4 considers other ResNets and the MiniONN architecture following COINN. Finally, in Section 6.5, we evaluate ImageNet against Cheetah, COINN, and CryptGPU.
@@
We plot a summary of the accuracy and inference time for CIFAR-100 in Figure 6. For both datasets (recall Figure 1 ), we observe that our work always gives the solution with the fastest inference time by a statistically significant amount. In terms of accuracy, our work is competitive with the state-of-the-art, but CryptGPU is always the most accurate as it can infer unmodified plaintext models. Our solution is faster than CryptGPU by 4×4$\times$4 ×, GForce by 5×5$\times$5 ×, and COINN by 18×18$\times$18 × on average in wide area networks. Our accuracies are competitive with state-of-the-art and plaintext solutions and stay stable (no escaping activations) with models containing up to $110$ layers and $23$ million parameters.
@@ 6.1 Experimental Setup
We develop an experimental setup that follows as closely as possible to the works we compare to [18, 31, 35]. We use CIFAR-10/100 [23] and ImageNet [10], the same common benchmark datasets as related work. Our model architectures include: MiniONN [26], VGG [34], and ResNets [16]. This covers models of depth $7$ to $110$ layers with the number of trainable parameters ranging from $0.2$ to $23$ million.
@@
All experiments are run on a machine with 32 CPU cores @ 3.7 GHz and 1 TB of RAM with two NVIDIA A100 with 80 GB of memory. We simulate network delay by calling the sleep function for the appropriate time whenever the client and server communicate. We simulate the LAN with 0.25 ms roundtrip delay and the WAN with 100 ms, following COINN [18]. We additionally evaluate a real WAN using AWS instances in Section 6.6. All experiments (except ImageNet) are repeated over multiple random seeds, and we report the mean and 95% confidence interval as shaded areas.
For all related work, we run our own benchmarks of their code unmodified. While results for ResNet32 and MiniONN appear in the Cheetah paper, there was no source code for these models so we only evaluate Cheetah on ImageNet. To use CryptGPU in practice, one must first train a model in PyTorch. For ImageNet, PyTorch provides pre-trained models. However, we will need to train a model for all other architectures and datasets. We simply use the same configurations as our PolyRelu models but with standard ReLUs. We include all source code to reproduce our results  [1].
@@
We introduce five new hyperparameters associated with our techniques: polynomial degree ($n$), polynomial approximation range ($\lambda$), polynomial regularization range ($\lambda_{reg}$), polynomial regularization coefficient ($\beta$), and polynomial regularization exponent ($\gamma$). The default values for each are decided by extensive grid searches. These parameters primarily affect the accuracy and not the inference time, except for the polynomial degree ($n$), which has a minor effect on the communication size (but not the rounds). We found the optimal quantization-aware polynomial degree ($n$) to be 4. We found that higher degrees than four increase the failure probability of truncation with minimal accuracy gains (Section 4.3). Conversely, lower degrees give lower accuracy (as shown in Figure 5). The polynomial used in all evaluations is: $0.31445312+0.5x+0.15625x^{2}-0.00292969x^{4}$.
For all models and datasets, we found a value of $\gamma=10$ performs well as it introduces a strong enough incentive for PolyReLU inputs to stay in range (smaller values did not) while keeping penalization for values in range practically 00 (if an input $x$ is within range, then $\frac{x}{\lambda_{reg}}<1\Rightarrow(\frac{x}{(\lambda_{reg})^{10}}\sim 0$). Larger values than this often lead to an infinite penalty term. We found that a polynomial approximation range $\lambda=5$ provides a good compromise between regularization and quantization. We found that smaller values of $\lambda$ destroy the accuracy during training, and larger values use too much precision, increasing the failure probability. The optimal polynomial regularization range ($\lambda_{reg}$) and polynomial regularization coefficient ($\beta$) vary per model and dataset, although we found $\lambda_{reg}=4.8$ (slightly tighter than $\lambda=5$) and $\beta=5\times 10^{-5}$ to be good default values. We use $p=10$ during ML training, but revert to CrypTen’s default precision of $p=16$ during inference in a $64$-bit ring.
We used Stochastic Gradient Decent as the optimizer with a learning rate of $0.013$ as the default. This includes a Cosine Annealing Learning Rate Scheduler with Linear Learning Rate Warmup of 5 epochs and decay $0.01$. We use a weight decay of either $10^{-4}$ or $5\times 10^{-4}$ and a momentum of $0.9$. We used a default batch size of $128$ and set the default number of Epochs to $185$. For some models, we tuned the learning rate, number of epochs, and regularization coefficient to achieve a slightly higher accuracy. We detail hyperparameters in our source code repository.
@@ 6.3 VGG-16 Architecture
We recall that we swap the MaxPool layers for AvgPool layers in the inference time evaluation. This comes at a cost to accuracy for the VGG architecture. Thus, to give the best scenario possible for GForce, we consider the accuracy of GForce with MaxPools and our work with AvgPool. We give the results in Table 3. As expected, GForce outperforms our work in accuracy (due to the MaxPools); however, only by a few percentage points. We train CryptGPU to use AvgPool and find it also loses a few percentage points, confirming our hypothesis that a MaxPool is necessary for high accuracy in VGG-16. We emphasize that our ResNet-18 result outperforms GForce’s VGG result in inference time and accuracy. Furthermore, ResNets are a more popular and compact architecture due to skip-connections.
@@ 6.4 Other Architectures
While GForce is the current state-of-the-art, COINN is a competitive solution that evaluates ResNet architectures. Thus, we also evaluate the same configurations as COINN. This includes the smaller MiniONN architecture, a ResNet-32, and a ResNet-110. We exclude GForce from this evaluation as they only evaluate VGG-16 models.
@@
We again swap all MaxPool layers for AvgPool in our work, and CryptGPU but leave COINN unmodified. We give the results in Figure 9. We observe that, over each of the three increasingly large architectures, the trends are similar and proportional to the number of parameters (0.2, 0.5, and 1.7 million parameters for MiniONN, ResNet-32, and ResNet-110, respectively). Across all architectures and network delays, our work outperforms COINN by a statistically significant amount (18×18$\times$18 × on average in WAN). We once again outperform CryptGPU in all evaluations [^1] All of which are statistically significant except MiniONN in LAN where the confidence intervals overlap slightly. with a 4×4$\times$4 × speed up on average in the WAN.
@@
CIFAR-10 / MiniONN PILLAR 88.1 $\pm$ 0.26 87.9 $\pm$ 0.46
@@
CIFAR-10 / ResNet-110 PILLAR 91.4 $\pm$ 0.18 91.4 $\pm$ 0.25
@@
CIFAR-100 / ResNet-32 PILLAR 67.8 $\pm$ 0.32 67.8 $\pm$ 0.47